OpenStudy (anonymous):

an air traffic controller is plotting the course of 2 jets scheduled to plan in 30 minutes. one aircraft is following a path defined by the equation 2x-y=9. the other by the equation 4x=-3y+13. Should the controller alter the paths of either aircraft?

OpenStudy (anonymous):

not sure what "scheduled to plan" means :/

OpenStudy (anonymous):

pellet sorry scheduled to land****

OpenStudy (anonymous):

shoot sorry***

OpenStudy (anonymous):

pellet? also, are x and y both land coordinates? or does x represent time and y distance?

OpenStudy (anonymous):

i said a swear word and it substituted that in for me :p and i have no clue the wording for these questions makes no sense

OpenStudy (anonymous):

oh awesome ok well if x and y are both land coordinates then we're both fudged. because i'm fairly sure there's nothing to hint at parametric equations and that would be the only way to solve it, if x is time and y is distance, then you want to solve the system of equations

OpenStudy (anonymous):

say x is time in hours, btw can you solve the system by yourself?

OpenStudy (anonymous):

i could 2 months ago :p

OpenStudy (anonymous):

effffff fudge this question

OpenStudy (anonymous):

haha ok :) well start by writing then in Ax + By = C form don't fudge it just yet, ti's easy!

OpenStudy (anonymous):

2x-y=9 4x+3y=13

OpenStudy (anonymous):

you could do elimination

OpenStudy (anonymous):

that's right, and that's usually the easiest method so go ahead and do that

OpenStudy (anonymous):

so 4x-2y=18 -4x+3y=13 _____________ -5y=5

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

perfect, now findx cause that's what we're interested in

OpenStudy (anonymous):

4x=-3(-1)+13 4x=16 x=4

OpenStudy (anonymous):

(4,-1) is the solution

OpenStudy (anonymous):

ok so that's in hours. the planes land in .5 hours, right? so will they crash?

OpenStudy (anonymous):

i dont know where to go after finding that

OpenStudy (anonymous):



OpenStudy (anonymous):

thats all i get out of finding the solution

OpenStudy (anonymous):

if they follow their paths as planned, then they meet (read: crash) in four hours. but, they both land in half an hour. so they don't crash, and nobody needs to change plans does that make sense?

OpenStudy (anonymous):

ohhh yesss

OpenStudy (anonymous):

makes perfect sense :)

OpenStudy (anonymous):

thank you soooo muchhhh :))))))))

OpenStudy (anonymous):

of course!

OpenStudy (anonymous):

happy to help

OpenStudy (anonymous):

i understand the rest of my questions somewhat